OpenStudy (poetryprincess):

Help Please? Factor The Trinomial. x^2-12x+27

OpenStudy (arabpride):

Simplifying x2 + -12x + 27 = 0 Reorder the terms: 27 + -12x + x2 = 0 Solving 27 + -12x + x2 = 0 Solving for variable 'x'. Factor a trinomial. ( -1x + 3)( -1x + 9) = 0 ~~~~~~~~~~~~~~~~~~~~~~ Can't forget to site my source: http://www.algebrahelp.com/calculators/equation/factoring/calc.do?equation=x^2+-+12x+%2B+27+%3D+0 Good Luck c:

OpenStudy (poetryprincess):

Thanks..